A balloon whose volume is 770 m3 is to be filled with hydrogen at atmospheric pressure (1.01×105Pa).If the hydrogen is stored in cylinders with volumes of 1.75 m3 at a gauge pressure of 1.17×106 Pa , how many cylinders are required? Assume that the temperature of the hydrogen remains constant.
Answer #1

Solution,

(1.17 x 10^6) + (1.01 x 10^5) = 1.27 x 10^6 Pa

= (1.27 x 10^6)/(1.01 x 10^5) = 12.58

No. of cylinder = 770/(1.75 x 12.58) ~ 35

Comment in case any doubt please rate my answer ...